OpenStudy (anonymous):

2g + 14 + 5g = 8g - 3 + 8

OpenStudy (anonymous):

are you solving for g?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok the answer is g=-1

OpenStudy (anonymous):

combine like terms first \[2g+14+5g=8g-3+8\] \[7g+14=8g+5\]

OpenStudy (anonymous):

yes satellite73 knows

OpenStudy (anonymous):

then subtract 7g from both sides \[14=8g-7g+5\] \[14=g+5\]

OpenStudy (anonymous):

subtract 5 from both sides \[14-5=g\] \[9=g\]

OpenStudy (anonymous):

answer is 9.

OpenStudy (anonymous):

o my bad i did it the other way